OpenStudy (adunb8):

finding magnitude of force P if resultant of 3 force is vertical.

OpenStudy (adunb8):

i have no idea how the answer became to be that.. anyone help?

OpenStudy (anonymous):

First of all do you understand the question?

OpenStudy (adunb8):

the question itself is confusing =/

OpenStudy (anonymous):

attachment is something different and the question is different .??

OpenStudy (adunb8):

oh wait i put wrong attachment sorry

OpenStudy (anonymous):

LOL

OpenStudy (adunb8):

@andras i got the right one

OpenStudy (anonymous):

You even have the solution. What is the problem?

OpenStudy (adunb8):

i dont know how the answer came to be that i need like steps...

OpenStudy (anonymous):

First he calculates all the forces in the x direction. That is \[R _{x}\] It has 3 parts as there are 3 forces. P is easy as it has only x direction, so its x component is the same as itself. For the other 2 vectors you need to understand the magical world of trigonometry, do you need explanation for that?

OpenStudy (adunb8):

it would be great if i can get that magical world trig.

OpenStudy (anonymous):

OK lets try, it is a pit hard to draw nicely here

OpenStudy (anonymous):



OpenStudy (adunb8):

cool i understood this part!

OpenStudy (anonymous):

That x is the x component of that 200lb force. How can we find x? With sin function http://www.mathsisfun.com/algebra/trigonometry.html There you can read a bit more about sin/cos/tan in a triangle. Can you answer me this: In that drawing what would sin40 equal to?

OpenStudy (adunb8):

would it be 200lbsin40 = Fx (x component?)

OpenStudy (anonymous):

Yes that is right, it comes from this \[\sin 40 =\frac{ x }{ 200 }\]

OpenStudy (adunb8):

i see! how would i draw 400lb part? its confusing.

OpenStudy (anonymous):

So you found the x component of one of the vectors.

OpenStudy (anonymous):



OpenStudy (anonymous):

Just keep in mind that in this case I labeled the x component of that vector as y. AND it is pointing to the negative direction

OpenStudy (adunb8):

how did you know to draw it like that?? that is the most confusing part sorry any good tricks in understanding to draw those confusing drawings?

OpenStudy (anonymous):

I draw the vector first, than the angle. When you have these 2 you need to find another angle that is 90 degrees so that you can use trigonometry. \[\cos 40=\frac{ y }{ 400 }\]

OpenStudy (adunb8):

oh i see but on the answer it has 200sin40 that we found and - 400 cos 40 . is 400cos40 x component or is it y component because it has \[\sum_{}^{}Fx\]